# Not All Sub-Neptune Exoplanets Have Magma Oceans

### v1

@@ 3.2 Parameter Exploration
Figure 3: Location of the envelope-mantle boundary in pressure-temperature space as a function of parameters within the model grid. Each curve represents the locus of boundary locations across a different parameter range, with an upwards facing triangle representing the maximum value and a downward triangle the minimum. In some cases, the model radius exceeds 4 $R_{\oplus}$ before the maximum (or minimum) is reached when $\mu=2.35\,$g/mol, meaning the following models are not shown at that MMW: $T_{0}>500\,$K, $P_{\rm rc}<31.6\,$bar and $x_{\rm env}>3.5\%$. Additionally, we do not include the model with $\mu=18\,$g/mol and $T_{0}=300\,$K as this is impacted by water condensation. The phase diagram of MgSiO3 is shown to indicate the phase at the envelope-mantle boundary for each model in the grid, including the solid-liquid and bridgmanite-post-perovskite (Brg-PPv) boundaries. Different line colors represent the MMW of the model. Fiducial values for parameters which are not varied in a given plot are shown below the panels.
@@
Another takeaway from Figure 3 is that, due to the nature of the silicate phase diagram, there are two possible ways to suppress a magma ocean: either the boundary can be too cold, leading to the freezing of magma into bridgmanite, or it can be too high-pressure, leading to a phase transition between liquid silicate and solid post-perovskite. Many of the model parameters which affect the existence of a magma ocean do so primarily in one of these two ways: for example, $T_{0}$ and $P_{\rm rc}$ primarily control the temperature at the boundary, whereas $M_{p}$ and $x_{\rm env}$ strongly affect the pressure. This provides insight into how and why certain sub-Neptunes may not possess magma oceans.
@@
Even for planets with $P_{\rm rc}$, $M_{p}$, and $T_{0}$ which maximize the likelihood of magma oceans, some combinations of $x_{\rm env}$ and MMW may lead to a solid surface. For example, a planet with $T_{0}=900\,$K, $M_{p}=8M_{\oplus}$, and $P_{\rm rc}=1\,$bar would not have a magma ocean for $x_{\rm env}>40$% and $\mu>15\,$g/mol. This highlights that even young (low $P_{\rm rc}$), hot (high $T_{0}$) planets could lack magma oceans if their envelopes are heavy enough. This has implications for the evolution of sub-Neptunes; some of them could spend very little time with a magma ocean, which could dramatically impact atmospheric and structural development (e.g. Y. Chachan & D. J. Stevenson, 2018; C. Seo et al., 2024; F. E. Rigby et al., 2024).
The boundary lines in Figure 4 between a solid and liquid surface are typically monotonic, with $x_{\rm env}$ decreasing as MMW increases. A notable exception to this occurs at $P_{\rm rc}=100\,$bar and $T_{0}=700\,$K, where we see an S-shaped turnover in the boundary curve. This means that lower MMW, low $x_{\rm env}$ model planets in this regime also lack magma oceans. This arises because the temperature profiles for these models (i.e., high $P_{\rm rc}$ and low $T_{0}$) are simply too cold at their base to sustain a magma ocean, resulting in a solid bridgmanite surface. This contrasts from the other models shown in Figure 4, where the magma ocean is typically suppressed due to very high pressures at the envelope-mantle boundary and a corresponding phase transition to post-perovskite. The former behavior only occurs for planets with very thin envelopes, often resulting in planetary radii that are so small that they no longer correspond to sub-Neptunes at all but rather super-Earths ($R_{p}<1.75\,R_{\oplus}$). We can conclude that this low-temperature and lower-pressure mantle solidification mode may be unlikely to occur within the sub-Neptune regime of exoplanets. We note that in Figure 4 we omit the boundary for $P_{\rm rc}=100\,$bar, $T_{0}=500\,$K, since this PPTT profile crosses over the liquid-vapor boundary for H2O and is thus impacted by condensation.
@@ 3.3 Implications for JWST Sub-Neptune Observations
Our sample of planets (shown in Table 2) includes several sub-Neptunes with published JWST spectra: TOI-421 b (B. Davenport et al., 2025); TOI-270 d (B. Benneke et al., 2024; M. Holmberg & N. Madhusudhan, 2024); GJ 1214 b (E. M.-R. Kempton et al., 2023; P. Gao et al., 2023; E. Schlawin et al., 2024; K. Ohno et al., 2025); GJ 9827 d (C. Piaulet-Ghorayeb et al., 2024); K2-18 b (N. Madhusudhan et al., 2023, 2025; R. Hu et al., 2025); and TOI-836 c (N. L. Wallack et al., 2024). We include four additional JWST targets whose spectra have not yet been published to fill out our coverage of mass and temperature space: TOI-824 b, TOI-1231 b, HD 207496 b and HD 86226 c.
@@
With one exception, we find that lower-MMW atmospheres (i.e., MMW is below the transitional value) possess magma oceans, while atmospheres with a MMW above the transitional value host solid surfaces. This result is in agreement with the findings from our large model grid, shown in e.g. Figure 4. The exception to this trend is HD 86226 c, for which lower-MMW atmospheres do not have a magma ocean, and higher-MMW atmospheres do. This is because HD 86226 c is the hottest planet in our sample while also having a relatively high density, meaning that at low MMW, the envelope fraction required to explain its mass and radius becomes very small. Since the atmosphere is so thin in these cases, the pressure and temperature at the boundary remain low enough that the silicate layer remains in the solid phase. A similar effect is seen in the lower left panel of Figure 4, where the transitional MMW for a magma ocean begins to decrease as $x_{\rm env}$ decreases. Recent HST observations of HD 86226 c hint at a high-metallicity atmosphere (K. A. Kahle et al., 2025), but further observations using JWST will be required to break the degeneracy with high-altitude clouds.
@@
For planets with measured MMW values from JWST observations, our results help to inform the likelihood of a present-day magma ocean. The transitional MMW for TOI-421 b is much higher than the upper limit of $\sim 2.7\,$g/mol found by B. Davenport et al. (2025) for all values of $P_{\rm rc}$, indicating that this planet should still host a magma ocean. For TOI-270 d ($\mu=5.47^{+1.25}_{-1.4}\,$g/mol, B. Benneke et al., 2024), GJ 1214 b ($\mu\gtrsim 9.8\,$g/mol, E. M.-R. Kempton et al., 2023; P. Gao et al., 2023), GJ 9827 d ($\mu=18.02^{+0.20}_{-0.61}\,$g/mol, C. Piaulet-Ghorayeb et al., 2024) and TOI-836 c ($\mu\gtrsim 6\,$g/mol, N. L. Wallack et al., 2024), our transitional MMW values span the reported MMW measurements. This suggests that further observational refinement of the MMW is needed to determine the likelihood of a magma ocean. Alternatively, detailed evolutionary modeling could help to refine the range of possible $P_{\rm rc}$ values (see Section 4.1). We note that molecular abundances in the atmosphere of TOI-270 d have been found to be consistent with the result of magma-atmosphere interactions (M. C. Nixon et al., in review). A MMW value has not been reported for K2-18 b. The interpretation of spectroscopic observations of this planet has been the source of considerable debate (e.g., S. P. Schmidt et al., 2025; R. Luque et al., 2025; L. Welbanks et al., 2025). Furthermore, due to the possibility of water condensation and cold trapping, the measured composition of the upper atmosphere of this planet may not be representative of the entire gaseous envelope. In this case, other atmospheric indicators may be used to determine the role of magma-atmosphere interactions in shaping this planet (F. E. Rigby et al., 2024; O. Shorttle et al., 2024). Overall, given the significant number of sub-Neptunes observed to have high-MMW envelopes, we highlight that present-day magma oceans on sub-Neptunes may not be ubiquitous.
@@
GJ 9827 d $3.02^{+0.58}_{-0.57}$ $1.89^{+0.16}_{-0.14}$ 620 – (all liquid) 17.04 7.89 V. M. Passegger et al. (2024)
TOI-270 d $4.78\pm 0.43$ $2.19\pm 0.07$ 365 5.93 4.63 2.67 V. Van Eylen et al. (2021)
HD 207496 b $6.10\pm 1.60$ $2.25^{+0.12}_{-0.10}$ 743 – (all liquid) 16.39 8.55 S. C. C. Barros et al. (2023)
TOI-421 b $6.70\pm 0.60$ $2.64\pm 0.08$ 981 15.73 13.12 10.51 A. F. Krenn et al. (2024)
HD 86226 c $7.25^{+1.19}_{-1.12}$ $2.16\pm 0.08$ 1311 – (all liquid) 2.67* 3.32* J. Teske et al. (2020)
GJ 1214 b $8.41^{+0.36}_{-0.35}$ $2.73\pm 0.03$ 596 11.16 9.20 3.32 A. S. Mahajan et al. (2024)
K2-18 b $8.63\pm 1.35$ $2.61\pm 0.09$ 284 4.63 3.32 – (all solid) B. Benneke et al. (2019)
TOI-836 c $9.60^{+2.70}_{-2.50}$ $2.59\pm 0.09$ 665 13.12 10.51 5.28 F. Hawthorn et al. (2023)
TOI-1231 b $15.40\pm 3.30$ $65^{+0.16}_{-0.15} 33$0 3.97 2.67 2.67 J. A. Burt et al. (2021)
TOI-824 b $18.47^{+1.84}_{-1.88}$ $2.93^{+0.20}_{-0.19}$ 1253 13.12 9.85 7.89 J. A. Burt et al. (2020)
Table 2: Selection of 10 sub-Neptunes with current or upcoming JWST observations, including transitional values of MMW, above which the planet no longer hosts a magma ocean for a given PPTT profile. Three values of $P_{\rm rc}$ are considered for each planet, with $\mu_{X\,\rm{bar}}$ representing the transitional MMW for $P_{\rm rc}=X\,$bar. Dashes indicate models that remain liquid or solid across all conditions. For HD 86226 c, MMWs above the transitional value result in a magma ocean (these values are marked by an asterisk). MMW ($\mu$) is reported in units of g/mol.
@@ 4.1 Connection to planetary evolution
As mentioned in Section 2, the $P_{\rm rc}$ parameter serves as a proxy for planetary age, with older planets having a deeper radiative-convective boundary (and therefore cooling over time). Our research suggests that some sub-Neptunes could therefore lose their magma oceans with time. This is analogous to the crystallization of the Earth’s magma ocean, which has been extensively studied (e.g., Y. Abe, 1993; C. E. Boukaré et al., 2015; M. D. Ballmer et al., 2017; J. Korenaga, 2023). Additional research has investigated the process of magma ocean crystallization for terrestrial exoplanets (e.g., L. Schaefer & L. T. Elkins-Tanton, 2018; C.-É. Boukaré et al., 2025). Given our findings, we suggest that further study into magma ocean crystallization for larger sub-Neptunes, where the conditions at the boundary are likely to be different (i.e., at higher pressures) to terrestrial exoplanets, is crucial to further our understanding of sub-Neptune evolution.
While $P_{\rm rc}$ serves as a useful approximate indicator of a planet’s age, formally connecting $P_{\rm rc}$ and age is non-trivial, and requires further assumptions about the planet’s formation conditions, the luminosity of the host star, and the cooling timescale of the planet (J. G. Rogers, 2025). Furthermore, the magma ocean crystallization process itself is highly complex (e.g., L. Schaefer & L. T. Elkins-Tanton, 2018) and would require much more sophisticated models than those used in this work. For these reasons, we chose to retain $P_{\rm rc}$ as a free parameter in the model, rather than tying our model to any specific assumptions about these factors. A future study could incorporate planetary evolution to determine the length of time for which a planet with certain properties would be expected to retain a magma ocean.
@@ 4.2 Model limitations and caveats
Understanding the interiors of large exoplanets requires improved data on material behavior at extreme pressures and temperatures. For instance, the EOS of water (H2O) at high $P$ and $T$ remains poorly constrained by experiments (B. Journaux et al., 2020). There is also some uncertainty in the silicate EOS and phase boundary locations, although the effect on mass-radius relations has been shown to be small for planets $2\,M_{\oplus}$ (C. Huang et al., 2022). For this initial parameter exploration, we have not considered the possibility of partial melting (G. Fiquet et al., 2010), instead assuming complete melting at the solid-liquid silicate phase boundary, an assumption which could be revisited in future work. Furthermore, while miscibility studies suggest that hydrogen/helium (H/He) and H2O mixtures approximate ideal mixing under certain conditions (F. Soubiran & B. Militzer, 2015), there are also some regions of parameter space where demixing could be expected (A. Gupta et al., 2025; S. Howard et al., 2025). Although the majority of models in our grid fall into the regime where a mixed envelope is expected, further refining of this boundary would allow for a better understanding of the structure of cooler sub-Neptunes. Further work should also explore the possibility of H2O and H2 mixing with the core and mantle, which has been proposed for planets in the sub-Neptune mass range (H. Luo et al., 2024; E. D. Young et al., 2025).
Our models follow the assumption that the planet’s envelope consists entirely of H, He and H2O. Although other chemical species, such as CH4 and CO2, have been identified in sub-Neptune upper atmospheres (e.g., B. Benneke et al., 2024), we do not possess high-pressure EOS data for these species and therefore they cannot be included in our models at present. Theoretical and experimental work to determine the high-pressure EOS of additional carbon-, nitrogen-, and sulfur-bearing chemical species would be invaluable to enable more detailed modeling efforts of the internal structures of sub-Neptunes.
Additionally, we do not account for the effects of interactions between the magma ocean and the envelope, which are likely to affect the composition of both components (H. E. Schlichting & E. D. Young, 2022). This process could, for example, enrich atmospheric MMW (C. Seo et al., 2024) and limit the total radius of the planet (E. S. Kite et al., 2019). However, given the substantial uncertainties which remain surrounding magma ocean-envelope interactions, due in part to a lack of relevant experimental data (F. E. Rigby et al., 2024), we choose not to model this effect and instead allow the atmospheric composition to vary freely. Further work into refining magma ocean-envelope interaction models would allow for this effect to be incorporated more directly into internal structure models.

### v2

@@ 3.2 Parameter Exploration
Figure 3: Location of the envelope-mantle boundary in pressure-temperature space as a function of parameters within the model grid. Each curve represents the locus of boundary locations across a different parameter range, with an upwards facing triangle representing the maximum value and a downward triangle the minimum. In some cases, the model radius exceeds 4 $R_{\oplus}$ before the maximum (or minimum) is reached when $\mu=2.35\,$g/mol, meaning the following models are not shown at that MMW: $T_{0}>500\,$K, $P_{\rm rc}<31.6\,$bar and $x_{\rm env}>3.5\%$. Additionally, we do not include the model with $\mu=18\,$g/mol and $T_{0}=300\,$K as this is impacted by water condensation. The phase diagram of MgSiO3 is shown to indicate the phase at the envelope-mantle boundary for each model in the grid, including the solid-liquid and bridgmanite-postperovskite (Brg-PPv) boundaries. Different line colors represent the MMW of the model. Fiducial values for parameters which are not varied in a given plot are shown below the panels.
@@
Another takeaway from Figure 3 is that, due to the nature of the silicate phase diagram, there are two possible ways to suppress a magma ocean: either the boundary can be too cold, leading to the freezing of magma into bridgmanite, or it can be too high-pressure, leading to a phase transition between liquid silicate and solid postperovskite. Many of the model parameters which affect the existence of a magma ocean do so primarily in one of these two ways: for example, $T_{0}$ and $P_{\rm rc}$ primarily control the temperature at the boundary, whereas $M_{p}$ and $x_{\rm env}$ strongly affect the pressure. This provides insight into how and why certain sub-Neptunes may not possess magma oceans.
@@
Even for planets with $P_{\rm rc}$, $M_{p}$, and $T_{0}$ which maximize the likelihood of magma oceans, some combinations of $x_{\rm env}$ and MMW may lead to a solid surface. For example, a planet with $T_{0}=900\,$K, $M_{p}=8M_{\oplus}$, and $P_{\rm rc}=1\,$bar would not have a magma ocean for $x_{\rm env}>40$% and $\mu>15\,$g/mol. This highlights that even young (low $P_{\rm rc}$), hot (high $T_{0}$) planets could lack magma oceans if their envelopes are heavy enough. This has implications for the evolution of sub-Neptunes; some of them could spend very little time with a magma ocean, which could dramatically impact atmospheric and structural development (e.g. Chachan_2018; Seo_2024; Rigby_2024).
The boundary lines in Figure 4 between a solid and liquid surface are typically monotonic, with $x_{\rm env}$ decreasing as MMW increases. A notable exception to this occurs at $P_{\rm rc}=100\,$bar and $T_{0}=700\,$K, where we see an S-shaped turnover in the boundary curve. This means that lower MMW, low $x_{\rm env}$ model planets in this regime also lack magma oceans. This arises because the temperature profiles for these models (i.e., high $P_{\rm rc}$ and low $T_{0}$) are simply too cold at their base to sustain a magma ocean, resulting in a solid bridgmanite surface. This contrasts from the other models shown in Figure 4, where the magma ocean is typically suppressed due to very high pressures at the envelope-mantle boundary and a corresponding phase transition to postperovskite. The former behavior only occurs for planets with very thin envelopes, often resulting in planetary radii that are so small that they no longer correspond to sub-Neptunes at all but rather super-Earths ($R_{p}<1.75\,R_{\oplus}$). We can conclude that this low-temperature and lower-pressure mantle solidification mode may be unlikely to occur within the sub-Neptune regime of exoplanets. We note that in Figure 4 we omit the boundary for $P_{\rm rc}=100\,$bar, $T_{0}=500\,$K, since this PPTT profile crosses over the liquid-vapor boundary for H2O and is thus impacted by condensation.
@@ 3.3 Implications for JWST Sub-Neptune Observations
Our sample of planets (shown in Table 2) includes several sub-Neptunes with published JWST spectra: TOI-421 b (Davenport_2025); TOI-270 d (Benneke_2024; Holmberg_2024); GJ 1214 b (Kempton_2023; Gao_2023; Schlawin2024; Ohno2025); GJ 9827 d (Piaulet-Ghorayeb_2024); K2-18 b (Madhusudhan_2023; Madhusudhan2025; Hu2025); and TOI-836 c (Wallack2024). We include four additional JWST targets whose spectra have not yet been published to fill out our coverage of mass and temperature space: TOI-824 b, TOI-1231 b, HD 207496 b and HD 86226 c.
@@
With one exception, we find that lower-MMW atmospheres (i.e., MMW is below the transitional value) possess magma oceans, while atmospheres with a MMW above the transitional value host solid surfaces. This result is in agreement with the findings from our large model grid, shown in e.g. Figure 4. The exception to this trend is HD 86226 c, for which lower-MMW atmospheres do not have a magma ocean, and higher-MMW atmospheres do. This is because HD 86226 c is the hottest planet in our sample while also having a relatively high density, meaning that at low MMW, the envelope fraction required to explain its mass and radius becomes very small. Since the atmosphere is so thin in these cases, the pressure and temperature at the boundary remain low enough that the silicate layer remains in the solid phase. A similar effect is seen in the lower left panel of Figure 4, where the transitional MMW for a magma ocean begins to decrease as $x_{\rm env}$ decreases. Recent HST observations of HD 86226 c hint at a high-metallicity atmosphere (Kahle2025), but further observations using JWST will be required to break the degeneracy with high-altitude clouds.
@@
For planets with measured MMW values from JWST observations, our results help to inform the likelihood of a present-day magma ocean. The transitional MMW for TOI-421 b is much higher than the upper limit of $\sim 2.7\,$g/mol found by Davenport_2025 for all values of $P_{\rm rc}$, indicating that this planet should still host a magma ocean. For TOI-270 d ($\mu=5.47^{+1.25}_{-1.4}\,$g/mol, Benneke_2024), GJ 1214 b ($\mu\gtrsim 9.8\,$g/mol, Kempton_2023; Gao_2023), GJ 9827 d ($\mu=18.02^{+0.20}_{-0.61}\,$g/mol, Piaulet-Ghorayeb_2024) and TOI-836 c ($\mu\gtrsim 6\,$g/mol, Wallack2024), our transitional MMW values span the reported MMW measurements. This suggests that further observational refinement of the MMW is needed to determine the likelihood of a magma ocean. Alternatively, detailed evolutionary modeling could help to refine the range of possible $P_{\rm rc}$ values (see Section 4.1). We note that molecular abundances in the atmosphere of TOI-270 d have been found to be consistent with the result of magma-atmosphere interactions (Nixon2025). A MMW value has not been reported for K2-18 b. The interpretation of spectroscopic observations of this planet has been the source of considerable debate (e.g., Schmidt2025; Luque2025; Welbanks2025). Furthermore, due to the possibility of water condensation and cold trapping, the measured composition of the upper atmosphere of this planet may not be representative of the entire gaseous envelope. In this case, other atmospheric indicators may be used to determine the role of magma-atmosphere interactions in shaping this planet (Rigby_2024; Shorttle_2024). Overall, given the significant number of sub-Neptunes observed to have high-MMW envelopes, we highlight that present-day magma oceans on sub-Neptunes may not be ubiquitous.
@@
GJ 9827 d $3.02^{+0.58}_{-0.57}$ $1.89^{+0.16}_{-0.14}$ 620 (All liquid) 17.04 7.89 Passegger2024
TOI-270 d $4.78\pm 0.43$ $2.19\pm 0.07$ 365 5.93 4.63 2.67 VanEylen2021
HD 207496 b $6.10\pm 1.60$ $2.25^{+0.12}_{-0.10}$ 743 (All liquid) 16.39 8.55 Barros2023
TOI-421 b $6.70\pm 0.60$ $2.64\pm 0.08$ 981 15.73 13.12 10.51 Krenn2024
HD 86226 c $7.25^{+1.19}_{-1.12}$ $2.16\pm 0.08$ 1311 (All liquid) 2.67* 3.32* Teske2020
GJ 1214 b $8.41^{+0.36}_{-0.35}$ $2.73\pm 0.03$ 596 11.16 9.20 3.32 Mahajan2024
K2-18 b $8.63\pm 1.35$ $2.61\pm 0.09$ 284 4.63 3.32 (All solid) Benneke2019
TOI-836 c $9.60^{+2.70}_{-2.50}$ $2.59\pm 0.09$ 665 13.12 10.51 5.28 Hawthorn2023
TOI-1231 b $15.40\pm 3.30$ $65^{+0.16}_{-0.15} 33$0 3.97 2.67 2.67 Burt2021
TOI-824 b $18.47^{+1.84}_{-1.88}$ $2.93^{+0.20}_{-0.19}$ 1253 13.12 9.85 7.89 Burt2020
Table 2: Selection of 10 sub-Neptunes with current or upcoming JWST observations, including transitional values of MMW, above which the planet no longer hosts a magma ocean for a given PPTT profile. Three values of $P_{\rm rc}$ are considered for each planet, with $\mu_{X\,\rm{bar}}$ representing the transitional MMW for $P_{\rm rc}=X\,$bar. Models that remain liquid or solid across all conditions are indicated in parentheses. For HD 86226 c, MMWs above the transitional value result in a magma ocean (these values are marked by an asterisk). MMW ($\mu$) is reported in units of g/mol.
@@ 4.1 Connection to planetary evolution
As mentioned in Section 2, the $P_{\rm rc}$ parameter serves as a proxy for planetary age, with older planets having a deeper radiative-convective boundary (and therefore cooling over time). Our research suggests that some sub-Neptunes could therefore lose their magma oceans with time. This is analogous to the crystallization of the Earth’s magma ocean, which has been extensively studied (e.g., Abe1993; Boukare2015; Ballmer2017; Korenaga2023). Additional research has investigated the process of magma ocean crystallization for terrestrial exoplanets (e.g., Schaefer2018; Boukare2025). Given our findings, we suggest that further study into magma ocean crystallization for larger sub-Neptunes, where the conditions at the boundary are likely to be different (i.e., at higher pressures) to terrestrial exoplanets, is crucial to further our understanding of sub-Neptune evolution.
While $P_{\rm rc}$ serves as a useful approximate indicator of a planet’s age, formally connecting $P_{\rm rc}$ and age is non-trivial, and requires further assumptions about the planet’s formation conditions, the luminosity of the host star, and the cooling timescale of the planet (Rogers2025). Furthermore, the magma ocean crystallization process itself is highly complex (e.g., Schaefer2018) and would require much more sophisticated models than those used in this work. For these reasons, we chose to retain $P_{\rm rc}$ as a free parameter in the model, rather than tying our model to any specific assumptions about these factors. A future study could incorporate planetary evolution to determine the length of time for which a planet with certain properties would be expected to retain a magma ocean.
@@ 4.2 Model limitations and caveats
Understanding the interiors of large exoplanets requires improved data on material behavior at extreme pressures and temperatures. For instance, the EOS of water (H2O) at high $P$ and $T$ remains poorly constrained by experiments (Journaux_2020). There is also some uncertainty in the silicate EOS and phase boundary locations, although the effect on mass-radius relations has been shown to be small for planets $2\,M_{\oplus}$ (Huang_2022). For this initial parameter exploration, we have not considered the possibility of partial melting (Fiquet2010), instead assuming complete melting at the solid-liquid silicate phase boundary, an assumption which could be revisited in future work. Furthermore, while miscibility studies suggest that hydrogen/helium (H/He) and H2O mixtures approximate ideal mixing under certain conditions (Soubiran_2015), there are also some regions of parameter space where demixing could be expected (Gupta2025; Howard2025). Although the majority of models in our grid fall into the regime where a mixed envelope is expected, further refining of this boundary would allow for a better understanding of the structure of cooler sub-Neptunes. Further work should also explore the possibility of H2O and H2 mixing with the core and mantle, which has been proposed for planets in the sub-Neptune mass range (Luo2024; Young25).
Our models follow the assumption that the planet’s envelope consists entirely of H, He and H2O. Although other chemical species, such as CH4 and CO2, have been identified in sub-Neptune upper atmospheres (e.g., Benneke_2024), we do not possess high-pressure EOS data for these species and therefore they cannot be included in our models at present. Theoretical and experimental work to determine the high-pressure EOS of additional carbon-, nitrogen-, and sulfur-bearing chemical species would be invaluable to enable more detailed modeling efforts of the internal structures of sub-Neptunes.
Additionally, we do not account for the effects of interactions between the magma ocean and the envelope, which are likely to affect the composition of both components (Schlichting_2022). This process could, for example, enrich atmospheric MMW (Seo_2024) and limit the total radius of the planet (Kite_2019). However, given the substantial uncertainties which remain surrounding magma ocean-envelope interactions, due in part to a lack of relevant experimental data (Rigby_2024), we choose not to model this effect and instead allow the atmospheric composition to vary freely. Further work into refining magma ocean-envelope interaction models would allow for this effect to be incorporated more directly into internal structure models.